OpenStudy (anonymous):

WILL GIVE MEDAL Half the perimeter of a building is 100 ft. The length of the building is 20 ft. more than the width. Find the dimensions of the building. (What are the 2 equations and how do I solve it?) Plss help!

OpenStudy (anonymous):

100(2) = 200 = Perimeter. 20+w = length; w = width 20+w + 20+w + w + w = 200

OpenStudy (anonymous):

Add like terms.

OpenStudy (anonymous):

@Sadworld so the dimensions are only "length and width"?

OpenStudy (anonymous):

Yes, for the perimeter.

OpenStudy (anonymous):

Wait, misread.

OpenStudy (sleepyjess):

For perimeter, you have 2l+2w=P. In this case they are telling you that L=w+20. So, if you substitute those into the equation, now you have 2(w+20)+2(w). Now, distribute.

OpenStudy (sleepyjess):

Does that make sense @eunalexis ?

OpenStudy (anonymous):

w(l)(h) = Dimension

OpenStudy (anonymous):

But it's in prism form. Do i have to add height?

OpenStudy (sleepyjess):

No, perimeter is only around the building, width and length. This can give another example, http://www.algebra.com/algebra/homework/conversion/Reduction-of-unit-multipliers.faq.question.209776.html

OpenStudy (sleepyjess):

They don't do it fully, but they show how to set up the equation.

OpenStudy (anonymous):

Thank you! :D @sleepyjess

OpenStudy (sleepyjess):

No problem! Do you have to solve it completely?

OpenStudy (anonymous):

Yes @sleepyjess

OpenStudy (sleepyjess):

Ok, so now that we have the equation, we need to distribute. Can you do that?

OpenStudy (anonymous):

Yep! Thank you! @sleepyjess

OpenStudy (sleepyjess):

Would you like me to finish it with you or do you have it from here?

OpenStudy (anonymous):

I can do it from here :) @sleepyjess

OpenStudy (sleepyjess):

Ok :)